OpenStudy (bieber896):

HELP!! Solve. 3(h-4)=-1/2(24-6h)

OpenStudy (loinducoeur):

First, expand the brackets. Then collect like terms. You should get the final answer by either dividing or multiplying the co-efficient. Do you want me to show you how to do it as well?

OpenStudy (bieber896):

aren't the like terms 6h and h?

OpenStudy (anonymous):

In depth First you are going to use the distributive property. 3(h-4) 3 times h = 3h 3 times 4 = 12 That comes out to 3h-12 = 1/2 (24-6h) So instead of dealing with the fraction on the other side i went 1 divide 2 that comes to a decimal (better than the fraction) anyways, 1/2 comes out to 0.5. So 0.5 times 24 = 12 0.5 times 6h = 2h The equation is now 3h - 12 = 12 - 3h Please correct me if im wrong. Thank you.

OpenStudy (bieber896):

@JustAnotherKitten Thanks so much for your help! :)

OpenStudy (anonymous):

You're welcome. If you need help with anymore just let me know and I will help!

OpenStudy (loinducoeur):

Why has the 1/2 turned positive, when the original question is a -1/2?

OpenStudy (bieber896):

isn't it a minus 1/2?

OpenStudy (loinducoeur):

If it was a -1/2, then overall, there would be no answer, other than 0; and that doesn't seem right either.